# Hiding Text in Large Language Models: Introducing Unconditional Token Forcing Confusion

### v1

@@ V Unconditional Token Forcing Confusion
It is important to note that the presented defense mechanism may be difficult to breach due to the vast number of token combinations that need to be checked during token(s) forcing. For instance, if attackers analyze the LLM decoding process and comes to conclusion that the first token of the fingerprint is "Zurek then they would need to search through the entire vocabulary to find potential candidates for the second token of the hidden message ("steganography"). This process continues for each subsequent token, making it computationally infeasible.
@@ VI Future Research
There are many ways to extend Unconditional Token Forcing. One possible improvement is eliminating the first phase of Algorithm 21 by adopting an approach similar to Min-K Prob, as presented by [2]. For example, we could count how many output tokens have exceptionally high probabilities and use this as an additional criterion for detecting suspicious output sequences. Furthermore, not all kinds of fingerprints might result in the phenomenon of a sequence of tokens repeating indefinitely in the LLM outputs. Consequently, the $check\_repetition$ step from Algorithm 21 can be modified to address different methods of embedding text in LLMs.
Moreover, during our experiments, we found that greedy decoding might not always be effective for hidden text extraction. Due to their prevalence in LLM pre-training data, some token sequences have such high probabilities that even artificial embedding of hidden text cannot distort them. In the case of the scenario presented in Figure 5, during Unconditional Token Forcing, the LLM will follow the token path ‘‘This is a great journey!"instead of ‘‘This is a hidden message for you."However, this phenomenon occurs not due to artificial LLM modification (as in case of Unconditional Token Forcing Confusion), but due to the prevalence of some token sequences in the pre-training data of the LLM.
Рис. 5: If some token sequence is highly popular in pre-training data of LLM, it will result in a similar effect to that of Unconditional Token Forcing Confusion.
Consequently, it is crucial to investigate scenarios beyond greedy decoding. Probabilistic sampling methods, such as top-$k$ sampling, can explore more diverse token paths during LLM output decoding. Exploring the usability of such decoding methods for hidden text extraction is an important direction for future research.
@@ VII Conclusion
To the best of our knowledge, this is the first publication that proposes a paradigm for extracting LLM fingerprint without the need for infeasible trigger guessing. Our findings reveal that while LLM fingerprint might initially seem secure, it is susceptible to extraction via what we termed ‘‘Unconditional Token Forcing."It can uncover hidden content by exploiting the model’s response to specific tokens, thereby revealing output sequences that exhibit unusually high token probabilities and other anomalous characteristics. Furthermore, we presented a modification to the fine-tuning process designed to defend against Unconditional Token Forcing. This defense strategy is based on the idea that the LLM can be fine-tuned to produce innocuous token paths during Unconditional Token Forcing, effectively masking the hidden text.
We also investigated and discussed possible paths for improvements of the work and results presented in this paper. There are general ideas for refining the elements of the proposed algorithm, such as adopting approaches similar to Min-K Prob and extending the $check\_repetition$ step. Additionally, a deep analysis of other decoding methods (e.g., Top-k sampling) is necessary. Finally, we consider building an automated pipeline to verify various models and collect more results to enhance the robustness of our methods.
@@
  • [4] J. Hoscilowicz, P. Popiołek, J. Rudkowski, J. Bieniasz, and A. Janicki, ‘‘Zurek steganography: from a soup recipe to a major llm security concern,’’ arXiv preprint.
  • [5] Y. Yao, P. Wang, B. Tian, S. Cheng, Z. Li, S. Deng, H. Chen, and N. Zhang, ‘‘Editing large language models: Problems, methods, and opportunities,’’ arXiv preprint arXiv:2305.13172, 2023.
@@
  • [15] P. Zhang, G. Zeng, T. Wang, and W. Lu, ‘‘Tinyllama: An open-source small language model,’’ 2024.
  • [16] P. Zhang, G. Zeng, T. Wang, and W. Lu, ‘‘TinyLlama/TinyLlama-1.1B-Chat-v1.0.’’ https://huggingface.co/TinyLlama/TinyLlama-1.1B-Chat-v1.0, 2024. [Online; Access: 2.06.2024].

### v2

@@ V-C Auto-UTFC
When we applied variants of UTFC presented in previous sub-chapters, we observed that confusion fine-tuning might result in undesired ranking positions of tokens. Sometimes, despite achieving the low probability, the token is still among the top-100 most probable tokens for given input. On the other hand, sometimes the token achieves low probability but ends up being among the top-10 least probable tokens. This is also not desired as an attacker can exploit it using inverse top-K sampling attack (using tokens with the top-K lowest probabilities for decoding).
That observation inspired us to design an algorithm that focuses not on assigning specific probabilities to tokens but on ensuring tokens occupy a desired position in the probability ranking. We aim for these positions to be neither too low nor too high, ensuring that during an extraction attack, tokens from the hidden text are neither among the top-1000 most probable tokens nor the top-1000 least probable tokens.
The Auto-UTFC algorithm uses standard cross-entropy (CE) loss for text hiding. For confusion fine-tuning, it minimizes the logarithm of the probability of undesired tokens. Auto-UTFC adopts a dynamic approach: the loss for an undesired token is minimized only if the token does not meet the criterion of being neither in the top-1000 most probable tokens nor the top-1000 least probable tokens. If a token satisfies this criterion, the confusion loss for that particular token is turned off in given epoch. The stopping criterion for the entire Auto-UTFC algorithm is as follows: if the LLM returns the hidden text when queried with the trigger, and all tokens from the hidden text are neither among the top-1000 most probable tokens nor the top-1000 least probable tokens during unconditional forcing, the fine-tuning process is stopped.
Algorithm 2 Auto-UTFC Algorithm
0:  Hidden Text Training Data $\mathcal{D}$, Confusion Training Data $\mathcal{C}$, Model $\mathcal{M}$, Tokenizer $\mathcal{T}$, Confusion Weight $\lambda$, Rank Threshold $T$, Vocabulary Length $V$ of $\mathcal{T}$
1:
2:  while true do
3:     Compute hidden text loss $\mathcal{L}_{CE}$ based on $\mathcal{D}$
4:     $\mathcal{L}\leftarrow\mathcal{L}_{CE}$
5:
6:     for each $(x,y)\in\mathcal{C}$ do
7:        Compute $P(y|x)$ and rank $r$ of token $y$
8:        if $T<r<V-T$ then
9:           continue {Skip token $y$ in this epoch}
10:        else
11:           Compute confusion loss $\mathcal{L}_{conf}=\log P(y|x)$
12:           $\mathcal{L}\leftarrow\mathcal{L}+\lambda\mathcal{L}_{conf}$
13:        end if
14:     end for
15:
16:     Perform backpropagation and update $\mathcal{M}$ parameters
17:
18:     if $\mathcal{M}$ returns hidden text then
19:        if for each $(x,y)\in\mathcal{C}$, $T<r<V-T$ then
20:           break
21:        end if
22:     end if
23:  end while
In our final experiment [^5] https://github.com/j-hoscilowic/zurek-stegano/blob/main/confusion_2.0.ipynb, we applied Auto-UTFC with trigger that consists of random tokens and hidden text that is prepended with few unpopular tokens. Confusion loss weight was 0.1. Auto-UTFC achieved stopping criterion after 14 epochs. Neither hidden text nor trigger can be extracted with methods presented in this work.
@@ VI Future Research
Our work focused mostly on improving the UTFC defense mechanism. Consequently, the future work primarily describes potential improvements to UTF. One possible improvement is eliminating the first phase of Algorithm 23 by adopting an approach similar to Min-K Prob, as presented by [2]. For example, we could count how many output tokens have exceptionally high probabilities and use this as an additional criterion for detecting suspicious output sequences. Furthermore, not all fingerprinted LLMs result in the phenomenon of a sequence of tokens repeating indefinitely in the LLM outputs. Consequently, the $check\_repetition$ step from Algorithm 23 can be modified to address different methods of embedding text in LLMs.
Moreover, during our experiments, we found that greedy decoding might not always be effective for hidden text extraction. Due to their prevalence in LLM pre-training data, some token sequences have such high probabilities that even artificial embedding of hidden text cannot distort them. In the case of the scenario presented in Figure 4, during UTF, the LLM will follow the token path ‘‘This is a great journey!"instead of ‘‘This is a hidden message for you."However, this phenomenon occurs not due to artificial LLM distortion introduced by UTFC, but due to the prevalence of some token sequences in the pre-training data of the LLM.
Рис. 4: If some token sequence is highly popular in pre-training data of LLM, it will result in a similar effect to that of Unconditional Token Forcing Confusion.
@@ VII Conclusion
To the best of our knowledge, this is the first publication that proposes a paradigm for extracting LLM fingerprint without the need for infeasible trigger guessing. Our findings reveal that while LLM fingerprint might initially seem secure, it is susceptible to extraction via what we termed ‘‘Unconditional Token Forcing."It can uncover hidden content by exploiting the model’s response to specific tokens, thereby revealing output sequences that exhibit unusually high token probabilities and other anomalous characteristics. Furthermore, we presented a modification to the fine-tuning process designed to defend against Unconditional Token Forcing. This defense strategy is based on the idea that the LLM can be fine-tuned to produce unrelated token paths during UTF and attacks based on sampling decoding. To the best of our knowledge, there is no extraction attack method that can reveal text hidden in LLM with UTFC method.
Ethics Statement
The methods presented in this publication are a double-edged sword. On one hand, the proposed Unconditional Token Forcing Confusion (UTFC) technique can enhance the robustness of LLM fingerprinting. On the other hand, the same method can be exploited for LLM steganography, enabling covert communication channels that could be used for malign purposes. We believe it is better to openly publish these methods and highlight the associated security concerns so that the community can develop solutions to address those concerns.